# Harder-Narasimhan filtrations of decorated vector bundles

### v1

@@ 3.2. Canonical reductions in the sense of Atiyah and Bott
We now make the connection between Harder-Narasimhan and canonical reductions in the works of Atiyah and Bott in [atiyah-bott_yangmills] and Biswas and Holla in [biswas-holla_harder-narasimhanreductionofprincbund]. In this subsection, we focus and recall the approach of Atiyah and Bott in [atiyah-bott_yangmills], for which we first need a technical result from [atiyah-bott_yangmills, Lemma 10.2].
@@
Let $\xi$ be a principal-$G$-bundle on a compact Riemann surface $X$, where $G$ is complex and reductive. By making the required choices, we fix a special-orthogonal vector bundle structure $(\mathrm{ad}(\xi),\beta,\tau)$ through Lemma 3.7. We define canonical reductions of $\xi$ using the following lemma, using that $\mathrm{ad}(\xi)$ is a Lie algebra bundle with fibers isomorphic to $\mathrm{Lie}(G)$.
@@
For the special-orthogonal Harder-Narasimhan filtration of $(\mathrm{ad}(\xi),\beta,\tau)$ from Theorems 3.4 and 3.5:
@@
Since $\mathrm{ad}(\xi)_{t}^{\perp}$ is a parabolic Lie algebra subbundle, we can state the definition of a canonical reduction of a principal bundle, following Atiyah and Bott in [atiyah-bott_yangmills, Page 589].
@@
The filtration in this definition exists due to Lemma 3.7 and Theorems 3.4 and 3.5. Let us revisit Example 3.9 to see how canonical reductions are related to Harder-Narasimhan filtrations.
@@
This is an example of how Harder-Narasimhan filtrations of vector bundles induce canonical reductions of frame bundles. In [atiyah-bott_yangmills, Pages 589-590], it is explained that Harder-Narasimhan filtrations of vector bundles always arise from canonical reductions of frame bundles.
It is fair to ask whether canonical reductions always exist in the general setting of principal-$G$-bundles, where $G$ is complex and reductive.
@@
For a principal-$G$-bundle $\xi$, a canonical reduction $s^{*}\xi$ to a parabolic subgroup $P$ of $G$ exists.
@@
Since $\mathrm{ad}(\xi)_{t}^{\perp}$ is a parabolic Lie algebra subbundle of $\mathrm{ad}(\xi)$, due to Lemma 3.11, its fibers are isomorphic to a parabolic Lie subalgebra of $\mathrm{Lie}(G)$. Through the exponential map, the image of this subalgebra generates a parabolic subgroup $P$ of $G$, such that the fibers of $\mathrm{ad}(\xi)_{t}^{\perp}$ are isomorphic to $\mathrm{Lie}(P)$. We fix an embedding $\iota:G\hookrightarrow\mathbf{GL}(r,\mathbb{C})$ from [borel_linearalgebraicgroups, I. 1.10 Proposition], which induces an embedding $\mathbf{GL}(\mathrm{Lie}(G))\hookrightarrow\mathbf{GL}(\mathrm{Lie}(\mathbf{GL}(r,\mathbb{C})))\cong\mathbf{GL}(\mathbb{C}^{r^{2}})$, to view $\mathrm{ad}(\xi)$ as a vector bundle of rank $r^{2}$. By taking the frame bundle of $\mathrm{ad}(\xi)$, Lemma 1.3 recasts the subbundle $\mathrm{ad}(\xi)_{t}^{\perp}$ as a section $s^{\prime}:X\rightarrow\mathrm{Fr}(\mathrm{ad}(\xi))/P^{\prime}$, where $P^{\prime}$ is the standard maximal parabolic subgroup of $\mathbf{GL}(\mathbb{C}^{r^{2}})$ that stabilizes $D\iota(\mathrm{Lie}(P))\subseteq\mathbb{C}^{r^{2}}$, where $D\iota$ is the derivative of $\iota$. For every $x\in X$, the corresponding fiber $\mathrm{Fr}(\mathrm{ad}(\xi))_{x}/P^{\prime}$ is a homogeneous space isomorphic to $\mathbf{GL}(\mathbb{C}^{r^{2}})/P^{\prime}$, we fix such an isomorphism (which is not unique). Every element of $\mathbf{GL}(\mathbb{C}^{r^{2}})/P^{\prime}$ corresponds to a linear map $D\iota(\mathrm{Lie}(P))\rightarrow\mathbb{C}^{r^{2}}/D\iota(\mathrm{Lie}(P))$. Since $s^{\prime}$ is a reduction of $\mathrm{Fr}(\mathrm{ad}(\xi))$, it maps $x\in X$ to an element corresponding to a linear map $D\iota(\mathrm{Lie}(P))\rightarrow D\iota(\mathrm{Lie}(G))/D\iota(\mathrm{Lie}(P))$. Through the exponential map, the image of this linear map is an element of $G/P$. A choice of isomorphism $\mathrm{Fr}(\mathrm{ad}(\xi))_{x}/P^{\prime}\cong\mathbf{GL}(\mathbb{C}^{r^{2}})/P^{\prime}$ also determines an isomorphism $\xi_{x}/P\cong G/P$ through $\iota$, using that the kernel of $\mathrm{Ad}:G\rightarrow\mathbf{GL}(\mathrm{Lie}(G))$ lies inside $P$. Since the construction of an element in $G/P$ above is compatible with local trivializations and gluing, we have found a reduction $s:X\rightarrow\xi/P$. This reduction is canonical by construction. ∎
The next natural question to ask is whether canonical reductions are unique. The choices made in giving $\mathrm{ad}(\xi)$ a special-orthogonal vector bundle structure in Lemma 3.7, which also affect the construction of a canonical reduction in Lemma 3.14, could lead to different canonical reductions. For this, the following lemma from [atiyah-bott_yangmills, Proposition 10.4] is useful.
@@
For a canonical reduction $s^{*}\xi$ of a principal-$G$-bundle $\xi$, and a homomorphism of complex reductive groups $G\rightarrow G^{\prime}$ that maps $\mathbf{R}(G)$ into $\mathbf{R}(G^{\prime})$, then the extension $(s^{*}\xi)(G^{\prime})$ is a canonical reduction of $\xi(G^{\prime})$.
This lemma explains the phenomenon we saw in Remark 3.6, but also that with the appropriate extensions of structure groups, the adjoint bundles $\mathrm{ad}(\xi)^{\perp}_{t}$ of canonical reductions are unique up to isomorphism. The uniqueness of canonical reductions then hinges upon how two reductions are related when their adjoint bundles are isomorphic, we claim that these reductions are conjugate to each other, a notion we now make precise. We recall the following result from [biswas-holla_harder-narasimhanreductionofprincbund, Lemma 2.2].
@@
Through $\mathrm{conj}_{g}$, we induce an extension $(s^{*}\xi)(P^{\prime})$ of $s^{*}\xi$ to $P^{\prime}$, and there exists an isomorphism $\psi:(s^{*}\xi)(P^{\prime})\rightarrow(s^{\prime*}\xi)$ of principal-$P^{\prime}$-bundles.
@@
$\mathrm{ad}(s^{*}\xi)_{x}=P\backslash(s(x)\times\mathrm{Lie}(P))\rightarrow\mathrm{ad}(s^{\prime*}\xi)_{x}=P^{\prime}\backslash(s^{\prime}(x)\times\mathrm{Lie}(P^{\prime})),\quad[v,Y]\mapsto[\eta_{x}(v),\mathrm{Ad}(g)(Y)],$
@@
Due to Lemma 3.16, two canonical reductions $s^{*}\xi$ and $s^{\prime*}\xi$ of $\xi$ to $P$ and $P^{\prime}$ are conjugate in the sense that there exists an isomorphism $\psi:(s^{*}\xi)(P^{\prime})\rightarrow(s^{\prime*}\xi)$ of principal-$P^{\prime}$-bundles. To recap, Harder-Narasimhan filtrations of vector bundles, as well as symplectic and special-orthogonal variants, generalize nicely to canonical reductions of principal-$G$-bundles, where $G$ is complex and reductive.
@@ 3.3. Canonical reductions in the sense of Biswas and Holla
The approach of Biswas and Holla in [biswas-holla_harder-narasimhanreductionofprincbund] to constructing canonical reductions is slightly different, using the notion of dominant characters of parabolic subgroups $P$ of $G$, as discussed first by Ramanathan in [ramanathan_stableprincipalbundles]. In that setting, a canonical reduction $s^{*}\xi$ of $\xi$ to a parabolic subgroup $P$ of $G$ has two key properties:
@@
Due to Lemma 3.16, and the fact that conditions (BH-i) and (BH-ii) are invariant under conjugation of $P$, we can assume without loss of generality that $P=P_{I}$ is a standard parabolic subgroup of $\mathbf{GL}(r,\mathbb{C})$. From [atiyah-bott_yangmills, Pages 589-590], we know that Harder-Narasimhan reductions of $E$ arise from canonical reductions $s^{*}\mathrm{Fr}(E)$ of $\mathrm{Fr}(E)$ to $P_{I}$. More precisely, by generalizing Lemma 1.3 from subbundles of $E$ to filtrations of $E$, $s^{*}\mathrm{Fr}(E)$ corresponds to a filtration of $E$, which is the Harder-Narasimhan filtration of $E$. To show (BH-i), we use that the extension $(s^{*}\mathrm{Fr}(E))(L_{I})$ to the Levi-factor $L_{I}$ of $P_{I}$ is the frame bundle of the product of the successive quotients $F_{m}$, $m=1,\ldots,t$ of the Harder-Narasimhan filtration of the vector bundle $E$ (the graded vector bundle associated to that filtration). Since the quotients are slope-semistable, their frame bundles are Ramanathan-semistable due to Theorem 1.11. Hence, $(s^{*}\mathrm{Fr}(E))(L_{I})$ is Ramanathan-semistable as a product of Ramanathan-semistable bundles. Now we prove (BH-ii). We recall that characters $\chi:P_{I}\rightarrow\mathbb{C}^{\times}$ are completely determined by their restriction $\chi|_{L_{I}}:L_{I}\rightarrow\mathbb{C}^{\times}$ to the Levi-factor, as in the Levi decomposition $P_{I}=U_{I}L_{I}$, $U_{I}$ is a unipotent group (and thus has trivial characters). Since $L_{I}$ is the product of groups $\mathbf{GL}(r_{m},\mathbb{C})$, $m=1\ldots,t$, the dominant characters $\chi:P_{I}\rightarrow\mathbb{C}^{\times}$ restricted to $L_{I}$ are easy to enumerate: They are generated by the characters $\chi_{m}$, $m=1\ldots,t-1$, that evaluate the determinant of the $m$-th diagonal block, and divide by the determinant of the $m+1$-st diagonal block, as seen in Example 3.19. Hence, it suffices to show the inequality in (BH-ii) for characters $\chi_{m}$, $m=1\ldots,t-1$. We have:
@@
To generalize Lemma 3.20 to the general principal-$G$-bundle case, where $G$ is complex reductive, and prove Theorem 3.17, we need to review some facts about characters. For the rest of this subsection, we fix a Cartan $T$ and a Borel $B$ of $G$, such that $T\subseteq B\subseteq G$ with the simple roots $\triangle$.
@@
For a principal-$G$-bundle $\xi$, and a standard parabolic subgroup $P_{I}$ of $G$:
@@
For a proof of this lemma, see the middle section of the proof of [*][Proposition 3.1]biswas-holla_harder-narasimhanreductionofprincbund. In the $\mathbf{GL}(r,\mathbb{C})$ case, we can think of $\chi_{\alpha}$, $\alpha\in I$ as corresponding to $\chi_{m}:P_{I}\rightarrow\mathbb{C}^{\times}$ from the proof of Lemma 3.20 and from Example 3.19. The above lemma tells us that it suffices to check the inequality of (BH-ii) on finitely many characters, as we did in the proof of Lemma 3.20. Another part of [biswas-holla_harder-narasimhanreductionofprincbund, Proposition 3.1] generalizes Step 1 in our Harder-Narasimhan filtration proofs from Theorems 3.1-3.5.
@@
Lemma 3.22 determines a reduction $s^{*}\xi$ of $\xi$, which we claim fulfills conditions (BH-i) and (BH-ii). The following lemma is taken from [*][Proposition 3.1]biswas-holla_harder-narasimhanreductionofprincbund.
@@
The existence of a reduction fulfilling (i) and (ii) follows from Lemma 3.22. Proving (BH-i) is done by first assuming $(s^{*}\xi)(L_{I})$ is not Ramanathan-semistable, and reaching a contradiction to $\deg(\mathrm{ad}(s^{*}\xi))=\deg_{\max}(\xi)$. For (BH-ii), we use Lemma 3.21 so that it suffices to show $\deg(\chi_{\alpha}(s^{*}\xi))>0$ for certain characters $\chi_{\alpha}$. For proof details, see [*][Proposition 3.1]biswas-holla_harder-narasimhanreductionofprincbund. Now that we have the existence of reductions fulfilling (BH-i) and (BH-ii), we know that they are unique up to conjugation. The following theorem is taken from [*][Theorem 4.1]biswas-holla_harder-narasimhanreductionofprincbund.
@@
Biswas and Holla provided a bundle-theoretic proof to the existence of canonical reductions of principal bundles. Earlier in [behrend_semi-stabilityofreductivegroupschemesovercurves], Behrend provides a proof of the same result using his theory of complementary polyhedra. From a principal-$G$-bundle $\xi$, he constructs its complementary polyhedron, a root-theoretic object with a notion of semistability equivalent to Ramanathan-semistability: condition (BH-i) corresponds to verifying the stability of a projection of the complementary polyhedron, and condition (BH-ii) is replaced with checking the positivity of the numerical invariants of the complementary polyhedron. Behrend’s approach to canonical reudctions is particularly useful when generalizing canonical reductions to other interesting objects, such as Higgs bundles and parahoric torsors.

### v2

@@ 4.2. Canonical reductions in the sense of Atiyah and Bott
We now make the connection between Harder-Narasimhan and canonical reductions in the works of Atiyah and Bott in [AB83] and Biswas and Holla in [BH04]. In this subsection, we focus and recall the approach of Atiyah and Bott in [AB83], for which we first need a technical result from [AB83, Lemma 10.2].
@@
Let $\xi$ be a principal $G$-bundle on a compact Riemann surface $X$, where $G$ is complex and reductive. By making the required choices, we fix a special-orthogonal vector bundle structure $(\mathrm{ad}(\xi),\beta,\tau)$ through Lemma 4.7. We define canonical reductions of $\xi$ using the following lemma, using that $\mathrm{ad}(\xi)$ is a Lie algebra bundle with fibers isomorphic to $\mathrm{Lie}(G)$.
@@
For the special-orthogonal Harder-Narasimhan filtration of $(\mathrm{ad}(\xi),\beta,\tau)$ from Theorems 4.4 and 4.5:
@@
Since $\mathrm{ad}(\xi)_{t}^{\perp}$ is a parabolic Lie algebra subbundle, we can state the definition of a canonical reduction of a principal bundle, following Atiyah and Bott in [AB83, Page 589].
@@
The filtration in this definition exists due to Lemma 4.7 and Theorems 4.4 and 4.5. Let us revisit Example 4.9 to see how canonical reductions are related to Harder-Narasimhan filtrations.
@@
This is an example of how Harder-Narasimhan filtrations of vector bundles induce canonical reductions of frame bundles. In [AB83, Pages 589-590], it is explained that Harder-Narasimhan filtrations of vector bundles always arise from canonical reductions of frame bundles.
It is fair to ask whether canonical reductions always exist in the general setting of principal $G$-bundles, where $G$ is complex and reductive.
@@
For a principal $G$-bundle $\xi$, a canonical reduction $s^{*}\xi$ to a parabolic subgroup $P$ of $G$ exists.
@@
Since $\mathrm{ad}(\xi)_{t}^{\perp}$ is a parabolic Lie algebra subbundle of $\mathrm{ad}(\xi)$, due to Lemma 4.11, its fibers are isomorphic to a parabolic Lie subalgebra of $\mathrm{Lie}(G)$. Through the exponential map, the image of this subalgebra generates a parabolic subgroup $P$ of $G$, such that the fibers of $\mathrm{ad}(\xi)_{t}^{\perp}$ are isomorphic to $\mathrm{Lie}(P)$. We fix an embedding $\iota:G\hookrightarrow\mathbf{GL}(r,\mathbb{C})$ from [Bor91, I. 1.10 Proposition], which induces an embedding $\mathbf{GL}(\mathrm{Lie}(G))\hookrightarrow\mathbf{GL}(\mathrm{Lie}(\mathbf{GL}(r,\mathbb{C})))\cong\mathbf{GL}(\mathbb{C}^{r^{2}})$, to view $\mathrm{ad}(\xi)$ as a vector bundle of rank $r^{2}$. By taking the frame bundle of $\mathrm{ad}(\xi)$, Lemma 2.3 recasts the subbundle $\mathrm{ad}(\xi)_{t}^{\perp}$ as a section $s^{\prime}:X\rightarrow\mathrm{Fr}(\mathrm{ad}(\xi))/P^{\prime}$, where $P^{\prime}$ is the standard maximal parabolic subgroup of $\mathbf{GL}(\mathbb{C}^{r^{2}})$ that stabilizes $D\iota(\mathrm{Lie}(P))\subseteq\mathbb{C}^{r^{2}}$, where $D\iota$ is the derivative of $\iota$. For every $x\in X$, the corresponding fiber $\mathrm{Fr}(\mathrm{ad}(\xi))_{x}/P^{\prime}$ is a homogeneous space isomorphic to $\mathbf{GL}(\mathbb{C}^{r^{2}})/P^{\prime}$, we fix such an isomorphism (which is not unique). Every element of $\mathbf{GL}(\mathbb{C}^{r^{2}})/P^{\prime}$ corresponds to a linear map $D\iota(\mathrm{Lie}(P))\rightarrow\mathbb{C}^{r^{2}}/D\iota(\mathrm{Lie}(P))$. Since $s^{\prime}$ is a reduction of $\mathrm{Fr}(\mathrm{ad}(\xi))$, it maps $x\in X$ to an element corresponding to a linear map $D\iota(\mathrm{Lie}(P))\rightarrow D\iota(\mathrm{Lie}(G))/D\iota(\mathrm{Lie}(P))$. Through the exponential map, the image of this linear map is an element of $G/P$. A choice of isomorphism $\mathrm{Fr}(\mathrm{ad}(\xi))_{x}/P^{\prime}\cong\mathbf{GL}(\mathbb{C}^{r^{2}})/P^{\prime}$ also determines an isomorphism $\xi_{x}/P\cong G/P$ through $\iota$, using that the kernel of $\mathrm{Ad}:G\rightarrow\mathbf{GL}(\mathrm{Lie}(G))$ lies inside $P$. Since the construction of an element in $G/P$ above is compatible with local trivializations and gluing, we have found a reduction $s:X\rightarrow\xi/P$. This reduction is canonical by construction. ∎
The next natural question to ask is whether canonical reductions are unique. The choices made in giving $\mathrm{ad}(\xi)$ a special-orthogonal vector bundle structure in Lemma 4.7, which also affect the construction of a canonical reduction in Lemma 4.14, could lead to different canonical reductions. For this, the following lemma from [AB83, Proposition 10.4] is useful.
@@
For a canonical reduction $s^{*}\xi$ of a principal $G$-bundle $\xi$, and a homomorphism of complex reductive groups $G\rightarrow G^{\prime}$ that maps $\mathbf{R}(G)$ into $\mathbf{R}(G^{\prime})$, then the extension $(s^{*}\xi)(G^{\prime})$ is a canonical reduction of $\xi(G^{\prime})$.
This lemma explains the phenomenon we saw in Remark 4.6, but also that with the appropriate extensions of structure groups, the adjoint bundles $\mathrm{ad}(\xi)^{\perp}_{t}$ of canonical reductions are unique up to isomorphism. The uniqueness of canonical reductions then hinges upon how two reductions are related when their adjoint bundles are isomorphic, we claim that these reductions are conjugate to each other, a notion we now make precise. We recall the following result from [BH04, Lemma 2.2].
@@
Through $\mathrm{conj}_{g}$, we induce an extension $(s^{*}\xi)(P^{\prime})$ of $s^{*}\xi$ to $P^{\prime}$, and there exists an isomorphism $\psi:(s^{*}\xi)(P^{\prime})\rightarrow(s^{\prime*}\xi)$ of principal $P^{\prime}$-bundles.
@@
$\mathrm{ad}(s^{*}\xi)_{x}=P\backslash(s(x)\times\mathrm{Lie}(P))$ $\rightarrow\mathrm{ad}(s^{\prime*}\xi)_{x}=P^{\prime}\backslash(s^{\prime}(x)\times\mathrm{Lie}(P^{\prime})),$
$[v,Y]$ $\mapsto[\eta_{x}(v),\mathrm{Ad}(g)(Y)],$
@@
Due to Lemma 4.16, two canonical reductions $s^{*}\xi$ and $s^{\prime*}\xi$ of $\xi$ to $P$ and $P^{\prime}$ are conjugate in the sense that there exists an isomorphism $\psi:(s^{*}\xi)(P^{\prime})\rightarrow(s^{\prime*}\xi)$ of principal $P^{\prime}$-bundles. To recap, Harder-Narasimhan filtrations of vector bundles, as well as symplectic and special-orthogonal variants, generalize nicely to canonical reductions of principal $G$-bundles, where $G$ is complex and reductive.
@@ 4.3. Canonical reductions in the sense of Biswas and Holla
The approach of Biswas and Holla in [BH04] to constructing canonical reductions is slightly different, using the notion of dominant characters of parabolic subgroups $P$ of $G$, as discussed first by Ramanathan in [Ram75]. In that setting, a canonical reduction $s^{*}\xi$ of $\xi$ to a parabolic subgroup $P$ of $G$ has two key properties:
@@
Due to Lemma 4.16, and the fact that conditions (BH-i) and (BH-ii) are invariant under conjugation of $P$, we can assume without loss of generality that $P=P_{I}$ is a standard parabolic subgroup of $\mathbf{GL}(r,\mathbb{C})$. From [AB83, Pages 589-590], we know that Harder-Narasimhan reductions of $E$ arise from canonical reductions $s^{*}\mathrm{Fr}(E)$ of $\mathrm{Fr}(E)$ to $P_{I}$. More precisely, by generalizing Lemma 2.3 from subbundles of $E$ to filtrations of $E$, $s^{*}\mathrm{Fr}(E)$ corresponds to a filtration of $E$, which is the Harder-Narasimhan filtration of $E$. To show (BH-i), we use that the extension $(s^{*}\mathrm{Fr}(E))(L_{I})$ to the Levi-factor $L_{I}$ of $P_{I}$ is the frame bundle of the product of the successive quotients $F_{m}$, $m=1,\ldots,t$ of the Harder-Narasimhan filtration of the vector bundle $E$ (the graded vector bundle associated to that filtration). Since the quotients are slope-semistable, their frame bundles are Ramanathan-semistable due to Theorem 2.11. Hence, $(s^{*}\mathrm{Fr}(E))(L_{I})$ is Ramanathan-semistable as a product of Ramanathan-semistable bundles. Now we prove (BH-ii). We recall that characters $\chi:P_{I}\rightarrow\mathbb{C}^{\times}$ are completely determined by their restriction $\chi|_{L_{I}}:L_{I}\rightarrow\mathbb{C}^{\times}$ to the Levi-factor, as in the Levi decomposition $P_{I}=U_{I}L_{I}$, $U_{I}$ is a unipotent group (and thus has trivial characters). Since $L_{I}$ is the product of groups $\mathbf{GL}(r_{m},\mathbb{C})$, $m=1\ldots,t$, the dominant characters $\chi:P_{I}\rightarrow\mathbb{C}^{\times}$ restricted to $L_{I}$ are easy to enumerate: They are generated by the characters $\chi_{m}$, $m=1\ldots,t-1$, that evaluate the determinant of the $m$-th diagonal block, and divide by the determinant of the $m+1$-st diagonal block, as seen in Example 4.19. Hence, it suffices to show the inequality in (BH-ii) for characters $\chi_{m}$, $m=1\ldots,t-1$. We have:
@@
To generalize Lemma 4.20 to the general principal $G$-bundle case, where $G$ is complex reductive, and prove Theorem 4.17, we need to review some facts about characters. For the rest of this subsection, we fix a Cartan $T$ and a Borel $B$ of $G$, such that $T\subseteq B\subseteq G$ with the simple roots $\triangle$.
@@
For a principal $G$-bundle $\xi$, and a standard parabolic subgroup $P_{I}$ of $G$:
@@
For a proof of this lemma, see the middle section of the proof of [BH04, Proposition 3.1]. In the $\mathbf{GL}(r,\mathbb{C})$ case, we can think of $\chi_{\alpha}$, $\alpha\in I$ as corresponding to $\chi_{m}:P_{I}\rightarrow\mathbb{C}^{\times}$ from the proof of Lemma 4.20 and from Example 4.19. The above lemma tells us that it suffices to check the inequality of (BH-ii) on finitely many characters, as we did in the proof of Lemma 4.20. Another part of [BH04, Proposition 3.1] generalizes Step 1 in our Harder-Narasimhan filtration proofs from Theorems 4.1-4.5.
@@
Lemma 4.22 determines a reduction $s^{*}\xi$ of $\xi$, which we claim fulfills conditions (BH-i) and (BH-ii). The following lemma is taken from [BH04, Proposition 3.1].
@@
The existence of a reduction fulfilling (i) and (ii) follows from Lemma 4.22. Proving (BH-i) is done by first assuming $(s^{*}\xi)(L_{I})$ is not Ramanathan-semistable, and reaching a contradiction to $\deg(\mathrm{ad}(s^{*}\xi))=\deg_{\max}(\xi)$. For (BH-ii), we use Lemma 4.21 so that it suffices to show $\deg(\chi_{\alpha}(s^{*}\xi))>0$ for certain characters $\chi_{\alpha}$. For proof details, see [BH04, Proposition 3.1]. Now that we have the existence of reductions fulfilling (BH-i) and (BH-ii), we know that they are unique up to conjugation. The following theorem is taken from [BH04, Theorem 4.1].
@@
Biswas and Holla provided a bundle-theoretic proof to the existence of canonical reductions of principal bundles. Earlier in [Beh95], Behrend provides a proof of the same result using his theory of complementary polyhedra. From a principal $G$-bundle $\xi$, he constructs its complementary polyhedron, a root-theoretic object with a notion of semistability equivalent to Ramanathan-semistability: condition (BH-i) corresponds to verifying the stability of a projection of the complementary polyhedron, and condition (BH-ii) is replaced with checking the positivity of the numerical invariants of the complementary polyhedron. Behrend’s approach to canonical reductions is particularly useful when generalizing canonical reductions to other interesting objects, such as Higgs bundles and parahoric torsors.